OpenStudy (anonymous):

Assuming that Caleb makes 70% of his free throws and shoots 10 free throws in a game, what is the probability that he will make? a) All 10 of his free throws in a game? b) Make seven or more of his free throws in a game? c) Find the mean number of free throws Caleb would make in the game.

OpenStudy (perl):

You can solve this using a binomial distribution

OpenStudy (anonymous):

how do i do that

OpenStudy (perl):

we know the probability of making any given free throw is .70

OpenStudy (anonymous):

yes

OpenStudy (perl):

lets define S = makes a free throw. F = fails to make a free throw

OpenStudy (perl):

S is short for success, F for fail

OpenStudy (anonymous):

okay

OpenStudy (perl):

a) we want the probability SSS SSS SSS SS , ten S's in a row each S has probability of 0.7 so you get 0.7 * 0.7 * 0.7 * ... *0.7 = 0.7^10

OpenStudy (anonymous):

so the answer to a is 0.0282475249

OpenStudy (perl):

correct

OpenStudy (anonymous):

okay so b?

OpenStudy (perl):

b) is a bit more tricky you want seven successful free throws for example: SSSSSSSFFFF

OpenStudy (anonymous):

seven or more

OpenStudy (perl):

right, so we also need eight S's, nine S's , and ten S's

OpenStudy (anonymous):

okay so we do .7^7 and .7^8 and .7^9 and .7^10?

OpenStudy (perl):

we also need probability of failures too

OpenStudy (perl):

for the case of exactly seven free throws SSS SSS S FFF .7^7 * .3^3 and you're not done, there are also more cases the order counts, you can have FFF SSS F SSSS

OpenStudy (anonymous):

0.0022235661

OpenStudy (perl):

so its pretty hard to do this by writing out all the possibilities. instead we use a formula, binomial formula

OpenStudy (perl):

Let's define a variable X = number of successful free throws (out of ten free throws)

OpenStudy (anonymous):

okay

OpenStudy (perl):

b) is asking P( X >= 7 ) = P( X = 7 ) + P( X= 8) + P( X = 9) + P( X = 10)

OpenStudy (anonymous):

okay

OpenStudy (perl):

So the general formula is P( X = k) = nCk * p^k * (1-p)^(n-k) here n = number of trials p = probability of success

OpenStudy (anonymous):

what is C and k?

OpenStudy (perl):

k is what you want , here k is 7. C stands for choose

OpenStudy (perl):

for example P( X = 7) = (10 choose 7) * .7^7 * (.3)^(3)

OpenStudy (anonymous):

oh okay

OpenStudy (perl):

so P( X >= 7 ) = P( X = 7 ) + P( X= 8) + P( X = 9) + P( X = 10) = (10 choose 7) * .7^7 * (.3)^(3) + (10 choose 8 ) * .7^8 * (.3)^2 + (10 choose 9) * .7^9 *.3^1 + (10 choose 10)* (.7^10)(.3)^0

OpenStudy (anonymous):

oh okay i see so the answer to b is 0.6496107184

OpenStudy (nincompoop):

he left, what the freckle

OpenStudy (perl):

sometimes they come back, or they have technical issues

OpenStudy (anonymous):

I'm back sorry

OpenStudy (perl):

did that make sense so far?

OpenStudy (anonymous):

yes

OpenStudy (nincompoop):

mine doesn't die

OpenStudy (perl):

you have to multiply by the binomial coefficient because you have more cases

OpenStudy (perl):

nin, what browser youre using? im using mozilla

OpenStudy (anonymous):

so the answer is not 0.6496107184?

OpenStudy (perl):

thats correct

OpenStudy (anonymous):

what is the binomial coefficient

OpenStudy (perl):

that tells you how many ways you can rearrange S and F

OpenStudy (perl):

if you have seven S's and three F's, how many ways you can rearrange them I can't write them all out , since there are 120 ways , or 10 choose 7 ways

OpenStudy (anonymous):

imconfused again

OpenStudy (perl):

which part is confusing

OpenStudy (anonymous):

how to multiply the binomial coeffienct

OpenStudy (perl):

you have to find the binomial coefficient

OpenStudy (perl):

or evaluate it, i should say

OpenStudy (anonymous):

im m sorry i don't get it

OpenStudy (anonymous):

@perl